Find h, the x coordinate of the vertex of this parabola. Consider the function:
f(x)=xsquared +4x+1.

Answer by stanbon(48579)

Find h, the x coordinate of the vertex of this parabola. Consider the function:
f(x)=x^2 +4x+1.
Complete the square:
x^2 + 4x + ? = y-1 + ?
---
x^2 + 4x + 4 = y - 1 + 4
---
(x+2)^2 = y + 3
---
h = 2
=======================================
Cheers,
Stan H.